What amount of money invested at 15% interest can provide the following scholarships. $30,000 at the end of each year for 6 years, $40,000 for the next 6 years and $50,000 thereafter?
Answer #1

The present value of the future cash flow can be written as

PV = 30,000(P/A, 15%,6)+40,000(P/A, 15%,6)(P/F,15%,6)+ 50,000(P/A, 15%,0)(P/F, 15%, 12)​​​​​

1 - 1.15-6 1-1.15-6 - +40,000 x = 0.15 - 0.15 PV = 30.000 x 50,000 0.15 x 1.1512 x 1.15-6+

50,000 PV = 30,000x3.78448+40,000x3.78448 x 1.15-+- 0.15 x 1.1512

50.000 PV = 113, 534.48+151, 379.31 x 1.15-6+ 0.15 x 1.1512

PV = 113,534.48 + 65, 445.45 + 62,302.38

PV = $ 241, 282.32

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.